how do i solve these proportions?
5/8=x+2/12
3x/4=x+7/6

Answer by harpazo(655)

5/8 = (x+2)/12
Cross-multiply and then divide.
8(x+2) = 5(12)
8x + 16 = 60
8x = 60-12
8x = 48
x = 48/8
x = 6
Do the other question the same way.